OpenStudy (anonymous):

What are the solutions that satisfy the nonlinear system of equations below? Select all answers that apply. y=3x 3x^2+y^2=12 A. (-1, 3) B. (10, -3) C. (1, 3) D. (-1, -3) E. (-3, 1) F. (3, -1)

OpenStudy (bahrom7893):

y=3x, therefore 3x^2 + (3x)^2 = 12 3x^2+9x^2 = 12 12x^2 = 12 x^2 = 1

OpenStudy (bahrom7893):

What possible values can x have?

OpenStudy (anonymous):

so that would be C. 1,3? is that all that could be applied

OpenStudy (bahrom7893):

no x can also be -1, because (-1)^2 = 1

OpenStudy (anonymous):

im not sure what values x may have the question doesn't say anymore then what i posted up

OpenStudy (bahrom7893):

Select all answers that apply. x=-1, what's y?

OpenStudy (anonymous):

Oh okay so im looking at answers A and C,

OpenStudy (anonymous):

y is 3

OpenStudy (anonymous):

when \[x ^{2}=1\] \[x={\pm}1\] so y=3 when x=1 and y=-3 when x=-1

OpenStudy (bahrom7893):

No x=-1 y = 3x = 3*(-1) = -3

OpenStudy (anonymous):

so C and D are the answers

OpenStudy (anonymous):

thanks @aditya96 for clarifying that(: